# Scalable and robust regression models for continuous proportional data

### v1

@@ 5.2 Robustness and scalability under spatial regression models
For predictive assessments, we selected 55,215 lakes from the LakeCat dataset, focusing on those with surface areas greater than 40,000 m2. In the spatial model with cobit link (7), making probabilistic predictions for these 55,215 lakes using a traditional GP prior is computationally prohibitive. Thus, we employ a nearest neighbor Gaussian process (NNGP) (Datta et al.,, 2016) on the spatial random effect $u(\cdot)$. We use a prior and algorithm similar to that in Section 5.2 for spatial cobin and micobin regression, and Stan for spatial beta regression; see Appendix S.6 for details. We ran three chains for each model for a total of 6,000 iterations per chain, discarding the first 1,000 samples from each as burn-in. The convergence diagnosis did not indicate any problems.
Table 3: Comparison of MMI data analysis results under three different models. Variables whose 95% credible intervals do not include zero are highlighted in bold. Asterisks indicate variables for which the 95% credible interval does not contain zero after removing two influential observations (See Figure 4).
($n=949$) Beta regression Cobin regression Micobin regression
Variable Estimate 95% CI Estimate 95% CI Estimate 95% CI
Intercept -2.363 (-4.160, -0.553) -2.106 (-3.859, -0.345) -1.797 (-3.551, -0.085)
agkffact -2.586 (-5.584, 0.330) -2.888 (-5.714, -0.003) -3.457 (-6.113, -0.800)
bfi 0.343 (0.016, 0.672) 0.293 (-0.022, 0.614) 0.229 (-0.082, 0.548)
cbnf 0.165 (-0.081, 0.412) 0.182 (-0.055, 0.420) 0.191 (-0.035, 0.425)
conif 0.081 (-0.002, 0.164) 0.093 (0.011, 0.176) 0.123 (0.044, 0.203)
crophay -0.079 (-0.250, 0.091) -0.063 (-0.231, 0.106) -0.054 (-0.213, 0.105)
fert -0.073 (-0.310, 0.158) -0.092 (-0.323, 0.132) -0.082 (-0.300, 0.138)
manure -0.048 (-0.202, 0.102) -0.036 (-0.182, 0.115) -0.029 (-0.173, 0.118)
pestic97 -0.014 (-0.106, 0.075) -0.021 (-0.108, 0.067) -0.025 (-0.108, 0.059)
urbmdhi -0.181 (-0.288, -0.076) -0.170 (-0.273, -0.067) -0.142 (-0.243, -0.041)
agkffact, soil erodibility factor; bfi, base flow index; cbnf, cultivated biological N fixation; conif, coniferous forest cover; crophay, crop/hay land cover; fert, synthetic N fertilizer use; manure, manure application; pestic97, 1997 pesticide use; urbmdhi, medium/high-density urban land cover. All variables are $\log_{2}(x+1)$ transformed. See Table S.3 in Appendix S.6 for a detailed description.
@@
Figure 3: Predicted MMI at 55,215 lakes from the spatial micobin regression model in terms of $E\{Y(s^{*})\mid X(s^{*})\}$ at unsampled location $s^{*}$. (Left) Posterior predictive mean of $E\{Y(s^{*})\mid X(s^{*})\}$. (Right) Posterior predictive standard deviation of $E\{Y(s^{*})\mid X(s^{*})\}$.
The estimated fixed-effect coefficients are summarized in Table 3. Our results show that decreased biotic integrity (low MMI) in US lakes is associated with high soil erodibility, low base flow index, low coniferous forest cover, and high urban land cover in the lake watershed area. The signs are generally sensible, and the variables selected based on 95% credible intervals coincide between cobin and micobin, but differ with beta regression. In terms of computation, both spatial cobin and micobin regressions outperform spatial beta regression in terms of mESS of $\bm{\beta}$ per unit time by more than a factor of 20. Regarding WAIC (Gelman et al.,, 2014) conditional on random effects, the beta, cobin, and micobin regression models yield -1093.4, -1103.5, and -1119.3, respectively, suggesting that micobin regression achieves the best predictive accuracy among the three models. Figure 3 illustrates the predicted MMI at 55,215 lakes, with the right panel reflecting increased uncertainty in southern California and western Texas due to sparse MMI data.
We further analyze quantile residuals (Dunn and Smyth,, 1996) to assess goodness-of-fit, with the model parameters fixed at the posterior mean (posterior median of $\lambda$ for cobin). Figure 4 shows that beta and cobin regression exhibit a lack of fit in the left tail, while micobin regression captures the left tail accurately but exhibits a slight under-estimation of the right tail. Further investigation reveals that the two influential observations with the lowest quantile residuals are identical for all three models, with MMI values of 0.02 and 0.021. We repeated the analysis, removing these observations. After removal, the variables selected based on 95% credible intervals that did not overlap zero changed for beta regression, while the results for the cobin and micobin regression remained the same. This is consistent with the beta regression score function being unbounded in $y$, leading to brittle results. See Appendix S.6 for details. Overall, the results suggest that the cobin and micobin regressions are more robust to observations near the boundaries, exhibit significantly better scalability, and achieve better predictive performance.
Figure 4: Comparison of quantile residuals for goodness-of-fit assessment, along with two observations corresponding to the lowest quantiles. The red line corresponds to the $y=x$ line.
@@ 7 Discussion
Beyond generalized linear (mixed) models, we anticipate that cobin and micobin distributions can be naturally incorporated into a more diverse family of models with continuous proportional data, such as tree ensembles (Schmid et al.,, 2013) or deep generative models (Loaiza-Ganem and Cunningham,, 2019). Developments of scalable inference methods are key to enabling such extensions with massive amounts of data. Similar to how Pólya-Gamma augmentation is connected with variational inference for logistic models (Durante and Rigon,, 2019), we hope that the proposed Kolmogorov-Gamma augmentation provides insight for the future development of approximate Bayesian inference methods, in addition to facilitating inference using MCMC.
@@ Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), by the National Science Foundation (NSF IIS-2426762), and by the Office of Naval Research (N00014-21-1-2510).
Software
Code to reproduce the analyses is available at https://github.com/changwoo-lee/cobin-reproduce.
@@
Section S.1 contains proofs of statements presented in the main article. Section S.2 contains pseudocode of the Kolomogorov-Gamma $(1,c)$ distribution and additional sampling details. Section S.3 contains a detailed derivation of cobin as an exponential dispersion model, density and cumulative distribution functions of cobin and micobin, and a discussion on the link and variance functions. Section S.5 contains derivations of Gibbs samplers as well as an EM algorithm. Finally, Section S.6 gives further information about benthic macroinvertebrate multimetric index data, analysis settings, and additional results.
S.1 Proofs
S.1.1 Proof of Theorem 2
@@
$\int_{0}^{\infty}e^{-t\kappa}\frac{\{\sinh(c/2)\}^{b}}{(c/2)^{b}}e^{-c^{2}\kappa/2}{p_{\textsc{kg}}}(\kappa;b,0)\mathrm{d}\kappa$ =$=\frac{\{\sinh(c/2)\}^{b}}{(c/2)^{b}}\int e^{-(-0.5c^{2}+t)\kappa}{p_{\textsc{kg}}}(\kappa;b,0)d\kappa$
@@
$k(\bm{\Theta}_{2}\mid\bm{\Theta}^{*}_{2})=\sum_{\bm{\lambda}\in\{1,\dots,L\}^{n}}p(\psi\mid\bm{\lambda})\left\{\int_{(0,\infty)^{n}}p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})p(\bm{\kappa}\mid\bm{\beta}^{*},\bm{\lambda})\mathrm{d}\bm{\kappa}\right\}p(\bm{\lambda}\mid\bm{\beta}^{*},\psi^{*})$ (S.2)
We aim to establish a lower bound of the curly bracket term in (S.2) that is independent of $\bm{\beta}^{*}$. This is achieved through two steps.
@@
We have a lower bound of the curly bracket term in (S.2):
@@
×$\times\prod_{i=1}^{n}\mathrm{sinhc}\left(\frac{\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}}{2}\right)^{\lambda_{i}}\exp\left(-\frac{\left\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}+(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}\right\}\kappa_{i}}{2}\right){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)$
@@
$\int_{0}^{\infty}\exp\left(-\frac{\left\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}+(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}\right\}\kappa_{i}}{2}\right){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)\mathrm{d}\kappa_{i}$ =$=\left\{\mathrm{sinhc}\left(\frac{\sqrt{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}|^{2}+|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}|^{2}}}{2}\right)\right\}^{-\lambda_{i}}$
@@
×$\quad\times 2^{-\sum_{i}\lambda_{i}}e^{-\sum_{i=1}^{n}\lambda_{i}^{2}/4}\exp\left\{-\frac{1}{2}\left(\frac{\bm{\beta}^{\mathrm{\scriptscriptstyle T}}X^{\mathrm{\scriptscriptstyle T}}X\bm{\beta}}{2}\right)\right\}$
@@
×$\quad\times\exp\left(-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}(I_{p}-(\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}X\Sigma_{\beta}^{1/2}/2+I_{p})^{-1})\mathbf{s}\right)$
×$\quad\times(2\pi)^{-p/2}|\Sigma^{\star}|^{-1/2}\exp\left(-\frac{1}{2}(\bm{\beta}-\bm{\mu}^{\star})^{\mathrm{\scriptscriptstyle T}}(\Sigma^{\star})^{-1}(\bm{\beta}-\bm{\mu}^{\star})\right)$
@@
It is easy to see how the first sum terms simplify into the $a_{n}^{R}(x)$ terms given in the statement of Proposition 2. Obtaining the second representation requires re-indexing the sum so that the even parts correspond to the term with $4x^{5/2}$ in the denominator and the odd parts correspond to the term with $x^{3/2}$ in the denominator. After that, the form for the $a^{L}_{n}(x)$ terms given in Proposition 2 emerges.
S.1.4 Proof of Lemma 1
@@
[…]
@@
$12^{9/2}\pi^{5/2}\exp\left(-2\pi^{2}t+\frac{1}{8t}+\frac{5}{2}\log t\right)=$ 1 (S.13)
whose solution is $t^{*}\approx 0.050239$ which lies within the bounds specified by Lemma 1. Since the derivative of the LHS of (S.13) is always negative on $t>0$, $t^{*}$ is unique.
@@
$m$ $1$ $2$ $3$ $4$ $5$ $6$
Prob. $0.127$ $2.068\times 10^{-4}$ $2.226\times 10^{-7}$ $3.124\times 10^{-11}$ $5.894\times 10^{-16}$ $1.513\times 10^{-21}$
@@
S.2 Pseudocode for KG(1,c) sampler
Algorithm S.1 Sampling from $\text{KG}(1,c)$
2:$\tilde{A}^{L}\leftarrow(|c|+2)\exp(-|c|/2)\texttt{pGIG}(t\mid p=-3/2,a=c^{2},b=1/4)$triangleright proportional to $A^{L}(c,t)$
3:$\tilde{A}^{R}\leftarrow 4\pi^{2}\exp\{-(2\pi^{2}+c^{2}/2)t\}/(2\pi^{2}+c^{2}/2)$triangleright proportional to $A^{R}(c,t)$
4:repeat
5:     Generate $U,V\sim U(0,1)$
7:         $X\leftarrow t+E/(2\pi^{2}+c^{2}/2)$ where $E\sim\text{Exp}(1)$triangleright Truncated exponential
8:     else
9:         $\triangleright$ Truncated GIG
10:              $X\sim\mathrm{GIG}(p=-3/2,a=c^{2},b=1/4)$triangleright If $c=0$, $X\sim\mathrm{InvGamma}(3/2,1/8)$
11:         until $X<t$
12:     end if
13:     $S\leftarrow a_{0}(X)$, $Y\leftarrow VS$, $m\leftarrow 0$
14:     repeat
15:         $m\leftarrow m+1$
16:         if $m$ is odd then
17:              $S\leftarrow S-a_{m}(X)$; if $Y<S$, then return $m$
18:         else
19:              $S\leftarrow S+a_{m}(X)$; if $Y>S$, then break
20:         end if
21:     until FALSE
22:until FALSE
The Algorithm S.1 describes the pseudocode for the KG$(1,c)$ sampler using the alternating series method. For the part involving the sampling and evaluation of the c.d.f of the GIG distribution with half-integer parameter $p=-3/2$, we employ the result of Peña and Jauch, (2025). When $c=0$, sampling from $\mathrm{GIG}(-3/2,c^{2},1/4)$ in line 10 is replaced with $\mathrm{InvGamma}(3/2,1/8)$, and c.d.f. evaluation in line 2 is replaced with c.d.f. of $\mathrm{InvGamma}(3/2,1/8)$. For the choice of optimal $t^{*}=0.050239$, we have $\texttt{pGIG}(t^{*}\mid p=-3/2,a=0,b=1/4)=0.1735472$.

### v2

@@ 4.3 Robustness to the presence of outliers
We fit spatial beta and cobin regression models with the cobit link (8) to $\mathcal{D}^{\circ}$ and $\mathcal{D}$, and spatial beta rectangular and micobin regression models to all three datasets, since these models can accommodate boundary values. We selected 55,215 lakes from the LakeCat dataset for prediction, focusing on those with surface areas greater than 40,000 m2. Since making probabilistic predictions for these 55,215 lakes using a traditional GP prior is computationally prohibitive, we employ a nearest neighbor Gaussian process (NNGP) (Datta et al., , 2016) on the spatial random effect. Since MMI values are unavailable for unsampled lakes, we assess predictive performance using WAIC (Gelman et al., , 2014) and leave-one-out cross-validation estimate using Pareto smoothed importance sampling (PSIS-LOO) (Vehtari et al., , 2024), both conditional on random effects. We use a prior and algorithm similar to that in Section 4 for spatial cobin and micobin regression, and Stan for spatial beta and beta rectangular regression; see Supplementary Section S.6.1 for details. We ran three chains for each model for a total of 6,000 iterations per chain, discarding the first 1,000 samples from each as burn-in. The convergence diagnosis with Gelman-Rubin statistic is less than 1.02 for all model parameters, indicating satisfactory convergence.
Table 5: Estimated fixed-effect coefficients ($\hat{\bm{\beta}}^{\circ}$) from four spatial regression models fitted on $\mathcal{D}^{\circ}$ ($n=949$) and 95% credible intervals. Bold entries indicate non-intercept coefficients whose 95% credible intervals exclude zero in $\mathcal{D}^{\circ}$ ($n=949$); asterisks indicate the same on fits to $\mathcal{D}$ ($n=947$)
Beta Cobin Beta rectangular Micobin
(Intercept) -2.363 (-4.160, -0.553) -2.106 (-3.859, -0.345) -1.609 (-3.370, 0.149) -1.797 (-3.551, -0.085)
agkffact -2.586(-5.584, 0.330) -2.888 (-5.714, -0.003) -3.041 (-5.902, -0.202) -3.457 (-6.113, -0.800)
bfi 0.343 (0.016, 0.672) 0.293 (-0.022, 0.614) 0.206 (-0.110, 0.523) 0.229 (-0.082, 0.548)
cbnf 0.165 (-0.081, 0.412) 0.182 (-0.055, 0.420) 0.201 (-0.034, 0.438) 0.191 (-0.035, 0.425)
conif 0.081 (-0.002, 0.164) 0.093 (0.011, 0.176) 0.106 (0.025, 0.187) 0.124 (0.044, 0.203)
crophay -0.079 (-0.250, 0.091) -0.063 (-0.231, 0.106) -0.036 (-0.204, 0.128) -0.054 (-0.213, 0.105)
fert -0.073 (-0.310, 0.158) -0.092 (-0.323, 0.132) -0.128 (-0.356, 0.097) -0.082 (-0.300, 0.138)
manure -0.048 (-0.202, 0.102) -0.036 (-0.182, 0.115) -0.018 (-0.166, 0.132) -0.029 (-0.173, 0.118)
pestic97 -0.014 (-0.106, 0.075) -0.021 (-0.108, 0.067) -0.035 (-0.121, 0.053) -0.025 (-0.108, 0.059)
urbmdhi -0.181 (-0.288, -0.076) -0.170 (-0.273, -0.067) -0.166 (-0.271, -0.062) -0.142 (-0.243, -0.041)
[…]
$\|\hat{\bm{\beta}}^{\circ}-\hat{\bm{\beta}}\|_{2}$ 0.743 0.444 0.098 0.069
[…]
We summarize the posterior mean estimate of fixed-effect coefficients from all four models based on $\mathcal{D}^{\circ}$ in Table 5, where non-intercept coefficients with 95% credible intervals excluding zero are emphasized in bold. Supplementary Section S.6.3 provides corresponding estimates based on $\mathcal{D}^{\star}$ and $\mathcal{D}$, as well as the average slopes (based on $\mathcal{D}^{\circ}$) for better interpretation of the results. The estimated signs of coefficients are generally sensible, where decreased biotic integrity (low MMI) in US lakes is associated with high soil erodibility, low base flow index, low coniferous forest cover, and high urban land cover in the lake watershed area.
Table 5 also highlights coefficients with asterisks whose 95% credible intervals exclude zero when the model is fitted to the dataset $\mathcal{D}$ without potential outliers. Notably, the beta regression yields different conclusions between $\mathcal{D}^{\circ}$ and $\mathcal{D}$, whereas cobin, beta rectangular, and micobin regression results remain consistent. This pattern aligns with the previous outlier robustness simulation, where beta regression exhibited poor stability and empirical coverage of $\bm{\beta}$, whereas micobin regression shows the highest stability. In terms of computational efficiency, both spatial cobin and micobin regressions based on Kolmogorov-Gamma augmentation outperform spatial beta and beta rectangular regression, achieving approximately a 20-fold improvement in mESS of $\bm{\beta}$ per unit time.
In environmental applications, it is also important to evaluate predictive performance in the presence of extreme observations rather than discarding them as mere anomalies. Lakes with low MMI likely reflect genuine, localized ecological degradation, and it is important that models accurately learn from such cases to make reliable predictions, particularly because similarly degraded lakes likely exist among the unsampled lakes. From this perspective, spatial micobin regression achieves the strongest predictive performance on $\mathcal{D}^{\circ}$ ($n=949$) with the lowest PSIS-LOO and WAIC. When the two low MMI lakes are excluded, spatial cobin regression achieves the best predictive performance on $\mathcal{D}$ ($n=947$), as its lighter tails compared to micobin yield sharper probabilistic predictions in the absence of outliers. Figure 3 illustrates the predicted MMI at 55,215 lakes from micobin regression fit based on $\mathcal{D}^{\circ}$, with the right panel reflecting increased uncertainty in southern California and western Texas due to sparse MMI data. Supplementary Section S.6.3 presents corresponding predictions from the beta, cobin, and beta rectangular regression models, along with their differences from the micobin predictions. Overall, the results suggest that the cobin and micobin regressions are more robust to observations near the boundaries, exhibit significantly better scalability, and achieve better predictive performance.
@@
Figure 3: Predicted MMI at 55,215 lakes from the spatial micobin regression model based on $\mathcal{D}^{\circ}$ in terms of $E\{Y(s^{*})\mid X(s^{*})\}$ at unsampled location $s^{*}$. (Left) Posterior predictive mean of $E\{Y(s^{*})\mid X(s^{*})\}$. (Right) Posterior predictive standard deviation of $E\{Y(s^{*})\mid X(s^{*})\}$.
@@ 6 Discussion
While the proposed cobin and micobin models demonstrate strong robustness properties and introduce a new framework for modeling continuous proportional data, our further investigation also revealed challenges in posterior computation under certain scenarios. When data are highly concentrated near the boundaries of the unit interval, the Markov chain based on Kolmogorov-Gamma augmentation can exhibit poor mixing behavior, analogous to known issues with Pólya-Gamma augmentation under highly imbalanced binary data (Johndrow et al., , 2019). Detailed theoretical and empirical results illustrating this phenomenon are provided in Supplementary Section S.7. Developing improved strategies to address these challenges, potentially through calibrated data augmentation (Duan et al., , 2018) or parameter expansion techniques (Zens et al., , 2024), represents an important direction for future work.
Beyond generalized linear (mixed) models, we anticipate that cobin and micobin distributions can be naturally incorporated into a more diverse family of models with continuous proportional data, such as tree ensembles (Stoddard et al., , 2008) or deep generative models (Loaiza-Ganem and Cunningham, , 2019). Extensions to multivariate outcomes, such as analyzing compositional data that lie on the simplex, also represent a promising future research avenue. Developments of scalable inference methods are key to enabling such extensions in large-scale settings. Similar to how Pólya-Gamma augmentation is connected with variational inference for logistic models (Durante and Rigon, , 2019), we hope that the proposed Kolmogorov-Gamma augmentation also provides insight for the future development of approximate Bayesian inference methods, in addition to facilitating computation using MCMC.
Data availability statement
Code to reproduce the analyses: https://github.com/changwoo-lee/cobin-reproduce.
@@ Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), by the National Science Foundation (NSF IIS-2426762), and by the Office of Naval Research (N00014-24-1-2626).
@@
Section S.1 contains sampling details from the Kolomogorov-Gamma $(1,c)$ distribution and the pseudocode of the algorithm. Section S.2 contains proofs of statements presented in the main article. Section S.3 contains a detailed derivation of cobin as an exponential dispersion model, density and cumulative distribution functions of cobin and micobin, and a discussion on the link and variance functions. Section S.4 contains derivations of Gibbs samplers as well as an EM algorithm. Section S.5 contains details of the simulation settings and additional results. Finally, Section S.6 gives further information about benthic macroinvertebrate multimetric index data, analysis settings, and additional results.
S.1 Kolmogorov-Gamma sampler
S.1.1 Kolmogorov-Gamma density
Proposition S.1.
$a_{n}(x;c,t)=\begin{cases}\left\{\sinh(c/2)/(c/2)\right\}\exp(-c^{2}x/2)a_{n}^{L}(x),&0<x<t,\\ \left\{\sinh(c/2)/(c/2)\right\}\exp(-c^{2}x/2)a_{n}^{R}(x),&t\leq x,\end{cases}$ (S.1)
S.1.2 Sampling KG(1,c) using alternating series method
The alternating series method (Devroye, , 1986) is an effective sampling algorithm when the target density $p(x)$ is computationally expensive to evaluate but can be approximated from above and below by a sequence of envelope functions $\{S_{m}(x)\}_{m=0}^{\infty}$, satisfying $S_{0}(x)>S_{2}(x)>\dots>p(x)>\dots>S_{3}(x)>S_{1}(x)$. With such an envelope function in hand, exact sampling of $X\sim p$ can be achieved with the following steps: (1) draw $X\sim q$ from the proposal distribution $q$, (2) generate $U\sim\mathrm{Unif}(0,Mq(X))$ where $\|p/q\|_{\infty}\leq M$, (3) repeat until $U\leq S_{m}(X)$ for odd $m$ or $U>S_{m}(X)$ for even $m$, (4) accept $X$ if $m$ is odd, repeat from (1) again if $m$ is even.
Similar to Devroye, (1986) and Polson et al., (2013), our choice of envelope function for sampling $\mathrm{KG}(1,c)$ is $S_{m}(x)=\sum_{n=0}^{m}(-1)^{n}a_{n}(x;c,t)$. For this choice to be a valid envelope function, $a_{n}(x;c,t)$ must be monotonically decreasing in $n$ for any fixed $x>0$ and parameter $c$. The following Lemma S.1 shows the valid range of $t$, an intersection between the range of $x$ where $a_{n}^{L}(x)$ and $a_{n}^{R}(x)$ are both monotonically decreasing in $n$.
Lemma S.1.
For any fixed $x>0$ and $c\in\mathbb{R}$, the sequence $a_{n}(x;c,t)$ defined in (S.1) is monotonically decreasing in $n$ when $\log(2)/(3\pi^{2})<t<0.25$, where $\log(2)/(3\pi^{2})\approx 0.0234$.
In what follows, we assume that $t$ satisfies Lemma S.1; the optimal choice of $t$ will be discussed soon. Since $S_{0}(x)=a_{0}(x;c,t)>{p_{\textsc{kg}}}(x;1,c)$ for all $x$, a natural choice of proposal $q$ that ensures $\|{p_{\textsc{kg}}}(\cdot;1,c)/q(\cdot)\|_{\infty}\leq M$ for some $M$ is $q(x;c,t)=M^{-1}a_{0}(x;c,t)$. Plugging in $a_{0}^{L}(x)$ and $a_{0}^{R}(x)$ in (S.1), the proposal distribution with density $q(x;c,t)$ is
[…]
where $\mathrm{GIG}(p,a,b)1(0<X\leq t)$ is a generalized inverse Gaussian (GIG) distribution with density proportional to $x^{p-1}\exp\{-(ax+b/x)/2\}$ truncated to $(0,t)$, $\mathrm{Exp}(a)1(t\leq X)$ is an exponential distribution with rate $a$ truncated to $t,\infty)$, $A^{L}(c,t)=\int_{0}^{t}a_{0}(x;c,t)\mathrm{d}x$, and $A^{R}(c,t)=\int_{t}^{\infty}a_{0}(x;c,t)\mathrm{d}x$. Compared to the Pólya-Gamma case (Polson et al., , 2013), which involves inverse Gaussian and exponential distributions in its proposal, our proposal is slightly more complicated, involving GIG and exponential distributions. Peña and Jauch, (2025) obtain results on the exact evaluation of the cdf and the sampling of GIG random variables that have a half-integer parameter p=1.5p=-1.5. We use these results in calculating $A^{L}(c,t)$ and sampling truncated GIG random variables.
Our $\mathrm{KG}(1,c)$ sampling algorithm is highly efficient. We show this by first investigating how often a proposal $X\sim q$ from (S.2) is accepted, based on the expected number of outer loop iterations $M=\int_{0}^{\infty}a_{0}(x;c,t)\mathrm{d}x=A^{L}(c,t)+A^{R}(c,t)$, and then by illustrating that with very high probability, only few series terms $S_{m}(x)$ need to be computed in the inner loop in order to decide whether to accept or reject $X$.
Proposition S.2.
The following statements hold:
  1. 1.
  2. 2.
S.1.3 Pseudocode for KG(1,c) sampler
Algorithm S.1 Sampling from $\text{KG}(1,c)$
[…]
4:repeat
5:  Generate $U,V\sim U(0,1)$
7:   $X\leftarrow t+E/(2\pi^{2}+c^{2}/2)$ where $E\sim\text{Exp}(1)$ $\triangleright$ Truncated exponential
8:  else
[…]
10:     $X\sim\mathrm{GIG}(p=-3/2,a=c^{2},b=1/4)$ $\triangleright$ If $c=0$, $X\sim\mathrm{InvGamma}(3/2,1/8)$
11:   until $X<t$
12:  end if
13:  $S\leftarrow a_{0}(X)$, $Y\leftarrow VS$, $m\leftarrow 0$
14:  repeat
15:   $m\leftarrow m+1$
16:   if $m$ is odd then
17:     $S\leftarrow S-a_{m}(X)$; if $Y<S$, then return $m$
18:   else
19:     $S\leftarrow S+a_{m}(X)$; if $Y>S$, then break
20:   end if
21:  until FALSE
22:until FALSE
The Algorithm S.1 describes the pseudocode for the KG(1,c)(1,c) sampler using the alternating series method. For the part involving the sampling and evaluation of the c.d.f of the GIG distribution with half-integer parameter p=3/2p=-3/2, we employ the result of Peña and Jauch, (2025). When $c=0$, sampling from $\mathrm{GIG}(-3/2,c^{2},1/4)$ in line 10 is replaced with $\mathrm{InvGamma}(3/2,1/8)$, and c.d.f. evaluation in line 2 is replaced with c.d.f. of $\mathrm{InvGamma}(3/2,1/8)$. For the choice of optimal $t^{*}=0.050239$, we have $\texttt{pGIG}(t^{*}\mid p=-3/2,a=0,b=1/4)=0.1735472$.
S.2 Proofs
S.2.1 Proof of Theorem 1
Proof.
$E\{\exp(-\kappa t)\}=\prod_{k=1}^{\infty}E\left\{\exp\left(-\frac{\epsilon_{k}t}{2\pi^{2}k^{2}}\right)\right\}=\prod_{k=1}^{\infty}\left(1+\frac{t}{2\pi^{2}k^{2}}\right)^{-b}=\left[\frac{(t/2)^{1/2}}{\sinh\{(t/2)^{1/2}\}}\right]^{b},$ (S.3)
S.2.2 Proof of Theorem 2
@@
$\int_{0}^{\infty}e^{-t\kappa}\frac{\{\sinh(c/2)\}^{b}}{(c/2)^{b}}e^{-c^{2}\kappa/2}{p_{\textsc{kg}}}(\kappa;b,0)\mathrm{d}\kappa$ $=\frac{\{\sinh(c/2)\}^{b}}{(c/2)^{b}}\int e^{-(-0.5c^{2}+t)\kappa}{p_{\textsc{kg}}}(\kappa;b,0)d\kappa$
@@
$k(\bm{\Theta}_{2}\mid\bm{\Theta}^{*}_{2})=\sum_{\bm{\lambda}\in\{1,\dots,L\}^{n}}p(\psi\mid\bm{\lambda})k(\bm{\beta}\mid\bm{\beta}^{*};\bm{\lambda})p(\bm{\lambda}\mid\bm{\beta}^{*},\psi^{*}),$ (S.5)
where the intermediate transition density of $\bm{\beta}$, conditional on $\bm{\lambda}$, is
$k(\bm{\beta}\mid\bm{\beta}^{*};\bm{\lambda})=\int_{(0,\infty)^{n}}p(\bm{\beta}\mid\bm{\kappa},\bm{\lambda})p(\bm{\kappa}\mid\bm{\beta}^{*},\bm{\lambda})\mathrm{d}\bm{\kappa}$ (S.6)
We aim to establish a lower bound of $k(\bm{\beta}\mid\bm{\beta}^{*};\bm{\lambda})$, expression (S.6), that is independent of $\bm{\beta}^{*}$. This is achieved through two steps.
@@
We have a lower bound of $k(\bm{\beta}\mid\bm{\beta}^{*};\bm{\lambda})$, expression (S.6):
@@
$\times\prod_{i=1}^{n}\mathrm{sinhc}\left(\frac{\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}}{2}\right)^{\lambda_{i}}\exp\left(-\frac{\left\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}+(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}\right\}\kappa_{i}}{2}\right){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)$
@@
$\int_{0}^{\infty}\exp\left(-\frac{\left\{(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*})^{2}+(\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta})^{2}\right\}\kappa_{i}}{2}\right){p_{\textsc{kg}}}(\kappa_{i};\lambda_{i},0)\mathrm{d}\kappa_{i}$ $=\left\{\mathrm{sinhc}\left(\frac{\sqrt{|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}^{*}|^{2}+|\bm{x}_{i}^{\mathrm{\scriptscriptstyle T}}\bm{\beta}|^{2}}}{2}\right)\right\}^{-\lambda_{i}}$
@@
$\quad\times 2^{-\sum_{i}\lambda_{i}}e^{-\sum_{i=1}^{n}\lambda_{i}^{2}/4}\exp\left\{-\frac{1}{2}\left(\frac{\bm{\beta}^{\mathrm{\scriptscriptstyle T}}X^{\mathrm{\scriptscriptstyle T}}X\bm{\beta}}{2}\right)\right\}$
@@
$\quad\times\exp\left(-\frac{1}{2}\mathbf{s}^{\mathrm{\scriptscriptstyle T}}(I_{p}-(\Sigma_{\beta}^{1/2}X^{\mathrm{\scriptscriptstyle T}}X\Sigma_{\beta}^{1/2}/2+I_{p})^{-1})\mathbf{s}\right)$
$\quad\times(2\pi)^{-p/2}|\Sigma^{\star}|^{-1/2}\exp\left(-\frac{1}{2}(\bm{\beta}-\bm{\mu}^{\star})^{\mathrm{\scriptscriptstyle T}}(\Sigma^{\star})^{-1}(\bm{\beta}-\bm{\mu}^{\star})\right)$
@@
It is easy to see how the first sum terms simplify into the $a_{n}^{R}(x)$ terms given in the statement of Proposition S.1. Obtaining the second representation requires re-indexing the sum so that the even parts correspond to the term with $4x^{5/2}$ in the denominator and the odd parts correspond to the term with $x^{3/2}$ in the denominator. After that, the form for the $a^{L}_{n}(x)$ terms given in Proposition S.1 emerges.
S.2.5 Proof of Lemma S.1
@@
Part 1. We begin by recalling the definition of $A^{L}(c,t)=\int_{0}^{t}a_{0}(x;c,t)dx$ and $A^{R}(c,t)=\int_{t}^{\infty}a_{0}(x;c,t)dx$, where the left and right envelopes are, from Proposition S.1 and (S.1),
@@
$2^{9/2}\pi^{5/2}\exp\left(-2\pi^{2}t+\frac{1}{8t}+\frac{5}{2}\log t\right)=1$ (S.17)
whose solution is $t^{*}\approx 0.050239$ which lies within the bounds specified by Lemma S.1. Since the derivative of the LHS of (S.17) is always negative on $t>0$, $t^{*}$ is unique.
@@
$m$ 11 22 33 44 55 66
Prob. $0.127$ $068\times 10^{-4} 2.2$$226\times 10^{-7}$ $3.124\times 10^{-11}$ $5.894\times 10^{-16}$ $1.513\times 10^{-21}$